A- A businessperson took out a loan of $ 120,000 from the Bank at a simple interest rate of 5% on June 12, which is due on August 16.

a) Find the number of days of the loan from June 12 until August 16.

b) Using exact interest, find the interest amount (I=?).

B- A dealer borrowed $5000 on a 110-day 2% simple interest note. He paid $1000 toward the note on day 30. On day 80 he paid an additional $1000. Assume 360-day year, what is his ending balance due?

c- A Supplier borrowed $ 89,600. The loan was for 14 months at a simple interest rate of 3%. What is the interest and the maturity value?





Answer by Boreal(15235)

a is 65 days or 65/365 of a year (they want exact, and that is exact, unless one was taught to always use 360, in which case the answer will be slight different.
I=prt=120000*0.05*65/365=$1068.49 interest
after day 30 interest if 5000*0.02*30/360=$8.33 so at day 30 he owed $5008.33-$1000 or $4008.33
in 80 more days it is $4008.33*0.02*80/3260=$17.81 interest so loan is now $4026.14. Pays off $1000 and now owes $3026.14
that *0.02*250/360 (days left in year) is $42.03 interest, so what will be due is $3068.17.
I=$89600*(0.03)(14/12)=$3136 interest
maturity value is that and starting amount of $92,736